HUBBLE SPACE TELESCOPE DAILY REPORT #2103 PERIOD COVERED: 0000Z (UTC) 04/15/98 - 0000Z (UTC) 04/16/98 Daily Status Report as of 106/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 7355 (Continuous Ultraviolet Monitoring of NGC 3516) The Space Telescope Imaging Spectrograph (CCD) was used to make continuous ultraviolet observations of galaxy NGC-3516. Recent intensive IUE monitoring of NGC-3516 and other AGN reveal rapid aperiodic variations down to the shortest time scales yet probed {~80 min}, strongly correlated throughout the UV and X-ray with no detectable interband lag. A wavelength calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 1.2 Completed WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 5 Successful : 3 The re-acquisition at 106/0200Z failed after five periods of loss of lock. (This, and the subsequent failure, used the same guide star pair.) HSTAR 6437 was written. Later, HSTAR 6438 described another re-acquisition failure at 106/0331Z that occurred since the roll error was high due to the previous failure. After the last failure, an ARU/PRT was uplinked at 106/0514Z. The following re-acquisition succeeded with difficulty. Each of the failures affected the iteration of proposal 7355 that will be covered by tomorrow's report. 2.2 FHST Updates: Scheduled : 2 Successful : 2 2.3 Operations Notes: The correctable SSR EDAC counts were cleared three times using ROP SR-1. ROP NS-5 was used to clear the engineering data report. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH